OpenStudy (anonymous):

How do you know which equation to solve for y and x with the substitution method?

OpenStudy (anonymous):

@hartnn @RadEn @robtobey

OpenStudy (anonymous):

@thomaster @Compassionate @Taylor<3sRin

OpenStudy (anonymous):

For example you have these two equations: 7x - 2y = 4 and 3x + y = 11

OpenStudy (anonymous):

How do you know which one you solve for y?

OpenStudy (anonymous):

@LastDayWork @annas @undeadknight26

hartnn (hartnn):

it really depends on the 2 equation, since i see "3x + y = 11" we can say its easy to isolate 'y' rather than x , so solve for y first in this case if we had something like x+12y =33, then its easier to isolate x here, so solve for x and use that as substution

OpenStudy (anonymous):

Is that always the case?

OpenStudy (anonymous):

For example 7x - 2y = 4, how can you tell which one is easier to isolate?

OpenStudy (lastdaywork):

@asdlkjfhaskldjfh It's just a matter of choice..

OpenStudy (anonymous):

Okay, so there's no way to tell

OpenStudy (lastdaywork):

Dividing by 2 is easier than 7; so i'll go with y :)

OpenStudy (anonymous):

Oh okay, that's a good point! haha thanks :)

hartnn (hartnn):

right, whatever you feel easier, go with that :)

OpenStudy (anonymous):

Okay, I would give you a medal but I can't thanks anyway :)

OpenStudy (lastdaywork):

Keep coming to OS :)

hartnn (hartnn):

no problem, i am anyways not here for medals, if you understand thats my medal :)

undeadknight26 (undeadknight26):

What he always says but he appreciates the appreciation XD

OpenStudy (texaschic101):

you deserve a medal hartnn :)

hartnn (hartnn):

thanks a lot :D

OpenStudy (lastdaywork):

I hope the discussion is not over my comment..it was directed at the newcomer..